# Hundred-Kilobyte Lookup Tables for Efficient Single-Image Super-Resolution

### v1

@@ Abstract
Conventional super-resolution (SR) schemes make heavy use of convolutional neural networks (CNNs), which involve intensive multiply-accumulate (MAC) operations, and require specialized hardware such as graphics processing units. This contradicts the regime of edge AI that often runs on devices strained by power, computing, and storage resources. Such a challenge has motivated a series of lookup table (LUT)-based SR schemes that employ simple LUT readout and largely elude CNN computation. Nonetheless, the multi-megabyte LUTs in existing methods still prohibit on-chip storage and necessitate off-chip memory transport. This work tackles this storage hurdle and innovates hundred-kilobyte LUT (HKLUT) models amenable to on-chip cache. Utilizing an asymmetric two-branch multistage network coupled with a suite of specialized kernel patterns, HKLUT demonstrates an uncompromising performance and superior hardware efficiency over existing LUT schemes.
Introduction
Single image super-resolution (SISR) is a long-standing problem in computer vision that involves generating a high-resolution (HR) image from a low-resolution (LR) image. The goal is to recover high-frequency details that are lost in a downsampled image. Classical non-deep-learning methods can be divided into two main categories: interpolation and sparse coding. Interpolation approaches (Keys 1981; Kirkland and Kirkland 2010), such as bicubic interpolation, are simple and computationally efficient. They involve upsampling the LR image using a fixed kernel, e.g., a bicubic filter. However, they often suffer from limited generalizability and produce blurry images. Sparse-coding approaches (Timofte, De Smet, and Van Gool 2013, 2015), on the other hand, can achieve better results than the former but are computationally expensive and require hand-crafted features.
@@ Introduction
Like many other problems in computer vision, deep learning has revolutionized SISR. The mapping between LR and HR images can be learned directly from data, often resulting in super-resolved images of much better quality than classical methods. However, deep neural networks (DNNs) require a large number of floating-point operations, resulting in a high power budget. Recently, researchers have explored new schemes that combine lookup tables (LUT) with deep-learning SISR (Jo and Kim 2021; Li et al. 2022; Ma et al. 2022). Specifically, this approach pre-computes all possible input-output pairs of a deep SR network and stores them into a LUT for retrieval, thus skipping expensive computation during inference. As a result, LUT-based methods achieve inference speeds comparable to interpolation methods (Jo and Kim 2021), but produce results with a learning-based quality. Moreover, LUT-based approaches are agnostic to software and hardware platforms, making them much more versatile than DNNs which require dedicated software and hardware support. Nonetheless, recent LUT-based SR schemes (Li et al. 2022; Ma et al. 2022) mainly focus on improving performance by enlarging the receptive field (RF), while overlooking the memory constraints. This contradicts the trending edge AI that often runs on edge devices strained by limited resources, e.g., the popular Ultra96 FPGA board has $<1$MB on-chip memory. In fact, existing LUT-based schemes generally require multi-megabyte LUTs where off-chip storage is inevitable. In addition, existing LUT-based approaches may still require expensive interpolation and/or floating-point operations, resulting in higher peak memory and energy costs. This highlights the challenge as well as the need for significant LUT size reduction and better architecture design to address the storage and power issues, while upkeeping performance.
@@
To this end, this work systematically explores the relationship between the number of input pixels, RF, and SR performance. Our study leads to the design of a family of models called Hundred-Kilobyte Look-up Tables (HKLUTs) which offer a compelling balance between memory and performance. To eliminate the time-consuming and energy-intensive interpolation, a two-branch architecture as in SPLUT (Ma et al. 2022) is employed. Yet instead of its symmetric branches (Ma et al. 2022), we exploit the idea of effective receptive field (ERF) (Luo et al. 2016) in different branches and reveal a larger RF is only required for the most-significant-bit (MSB) branch representing contextual semantic information, whereas a much smaller RF is sufficient for processing the least-significant-bit (LSB) branch.
@@
This key insight leads to an innovative asymmetric parallel structure, which largely reduces the storage space by nearly half without sacrificing performance. Next, we extend our model to multistage by cascading LUTs to enlarge the RF. Our proposed architecture, in contrast to prior works, allows information exchange between two branches during stage transition, resulting in an uncompromising performance. Moreover, we are the first to utilize progressive upsampling on LUT-based approaches, which elegantly reduces the LUT size by half while even improving the final performance. Compared to other state-of-the-art (SOTA) LUT-based schemes, HKLUTs deliver competitive performance while being >10×>10$\times$> 10 × smaller in size (cf. Fig. 1). The main contributions of this paper are threefold:
@@
The quest for high-quality SR on edge devices has fueled the interest and research of effective and efficient SISR schemes in recent years, which can be broadly categorized into two types: classical and deep-learning methods.
@@
Classical Approaches. Basic linear interpolation methods (Keys 1981; Kirkland and Kirkland 2010), though simple and efficient, are limited in their abilities to adapt to the image content and often result in blurry images. Early example-based work (Timofte, De Smet, and Van Gool 2013, 2015) learned sparse dictionaries over pixels or patches to build a compact representation between LR and HR pairs. During inference, the HR output is obtained using the pre-computed projection matrix with the input features. With low computation requirements, RAISR (Romano, Isidoro, and Milanfar 2016) proposed an effective sharpening algorithm to enhance the visual quality of over-smoothed images caused by cheap interpolation methods. However, such sparse coding models remain time-consuming to compute the sparse representations, while RAISR still needs local gradient statistics and fails to provide optimal solutions due to its non-differentiable hashing operations.
@@
Deep Learning-based Approaches. Deep learning methods have shown promising results in reconstructing image details for accurate super-resolution. Many deep learning SR works have been developed to reduce computational burdens by carefully adjusting the network structures, including, ESPCN (Shi et al. 2016a), CARN (Ahn, Kang, and Sohn 2018), RRDB (Wang et al. 2018) and IMDN (Hui et al. 2019). FALSR (Chu et al. 2021) proposed using neural architecture search (NAS) for both micro (block structure) and macro (layer-wise connection) levels. LapSRN (Lai et al. 2017) and its extension MS-LapSRN (Lai et al. 2018) employ Laplacian Pyramid to generate multi-scale predictions and progressively reconstruct HR images in multiple steps. FMEN (Du et al. 2022) shined in the NTIRE 2022 challenge with its low memory cost and short runtime through the enhanced residual block and high-frequency attention module. Although some models can be executed in real-time on GPUs, it is impractical to deploy them on edge devices with constrained hardware resources.
@@
LUT-based Approaches. While deep learning-based methods can produce impressive results, they often require heavy computations and large storage. To accelerate the inference of DNNs, dedicated software and hardware platforms such as CUDA and GPUs are required. This poses challenges for cost-sensitive and resource-restrictive edge devices such as smartphones and TVs. To address this, SRLUT (Jo and Kim 2021) fuses DNN and LUT for SR applications by pre-computing and caching input-output pairs of a pre-trained SR network into a LUT for retrieval at test time. This approach eliminates expensive computation during DNN inference and achieves similar speed but better SR results than classical approaches. The number of input pixels is restricted to 4, preventing exponential growth in storage space for the resulting LUT. The rotation ensemble is proposed for both training and inference to improve SR performance. To further reduce the LUT size, input values are discretized into a fewer number of fixed levels, while interpolation is used to recover the quantization loss. Two independent works around the same time, namely, MuLUT and SPLUT (Li et al. 2022; Ma et al. 2022), propose to expand the RF covered by cascading LUTs. MuLUT proposes the use of multiple complementary LUTs together with the rotation ensemble to solve the intrinsic limitation of a small RF in SRLUT. It also designs a LUT-aware fine-tuning strategy to boost performance. However, MuLUT preserves the costly interpolation step of SRLUT. On the other hand, SPLUT (Ma et al. 2022) exploits a novel parallel framework where the quantization loss is compensated by using two parallel branches that separately handle the most significant 4 bits and least significant 4 bits. This eliminates the time-consuming interpolation step for a faster inference speed. SPLUT also suggests enlarging the RF through feature aggregation along the horizontal and vertical directions. Nonetheless, despite showing improved performance compared to SRLUT, both SPLUT and MuLUT require several times more storage space.
Development of HKLUT
Reduced Number of Input Pixels
@@
The storage size of a LUT can be calculated using the formula $v^{n}\times r^{2}$, where $v$, $n$, and $r$ are the number of quantization levels, number of input pixels, and upscaling factor, respectively. The first term $v^{n}$ determines the number of entries, whereas the second term $r^{2}$ refers to the length of the output vector of each LUT. It is obvious that reducing the number of input pixels $n$ can result in an exponential reduction in storage space. Previous studies have primarily focused on expanding the RF by utilizing a fixed number of pixels (say, 4) to improve the final performance. SPLUT (Ma et al. 2022) uses feature aggregation to expand the RF, namely, the output feature maps generated from the previous LUT are divided into groups along the channel dimension and then padded horizontally or vertically. The padded feature maps are then added to generate a single-channel feature map, which serves as the input for subsequent LUTs. However, in order to generate multi-channel feature maps, the size of intermediate LUTs grows linearly with the number of output channels. This goes against our primary goal, which is to reduce storage space. Subsequently, we deliberately omit feature aggregation. On the other hand, SRLUT (Jo and Kim 2021) and MuLUT (Li et al. 2022) employ rotation ensemble to increase the area being covered. The final prediction $\hat{y}_{i}$ is computed as follows:
@@
Figure 2: Three types of kernels covering a $3\times 3$ RF with rotation ensemble via different numbers of input pixels. (a) SRLUT (Jo and Kim 2021) (b) HDLUT (c) LLUT.
@@
Asymmetric Parallel Structure
Previous studies have attempted to reduce the size of LUTs by limiting the number of quantization levels to 17 and using interpolation to recover the quantization loss (Jo and Kim 2021; Li et al. 2022). However, the interpolation stage is time-consuming and energy-intensive, hindering the practicality on edge devices. To address this, SPLUT (Ma et al. 2022) proposes a parallel branch structure that avoids the need for interpolation. The structure divides the 8-bit input image into its most and least significant bits, viz. MSB and LSB branches each with 4 bits (viz. 16 levels) of information. The final SR image is obtained by fusing the results of the two branches at the last stage. While this approach gives promising results, it overlooks an important fact, namely, the effective receptive fields (ERFs) of MSB and LSB branches are completely different as they represent different kinds of information. Therefore, using the same sets of LUTs for both MSB and LSB branches might not be an optimal choice.
@@
Multistage
LUT-based methods (Ma et al. 2022; Li et al. 2022) have demonstrated the efficacy of cascaded stages to enlarge the RF for better performance. SPLUT (Ma et al. 2022) splits the image into the MSB and LSB branches and deals with them separately with uniform quantization until the final addition, which prevents each branch from getting information from the other. MuLUT (Li et al. 2022) utilizes quantization and interpolation to build the connection between stages, which requires later fine-tuning to compensate for the loss of information caused by quantization.
Here, we propose a new framework to construct the cascaded stages as illustrated in Fig. 5. In each stage, we split the 8-bit input image into MSB and LSB branches for different LUT blocks to extract the features and upscale the resolution, then combine the features from both branches to compose a new enlarged image. The input and output of each stage are 8-bit images with different resolutions if the upscaling factor is not equal to 1. Such a module can then be stacked to constitute cascaded stages, wherein the latter stage can attain information from the two branches of the former stages and residual connections can be conveniently implemented. The entire flow requires no interpolation, which effectively speedups inference.
@@
Nearest (Bevilacqua et al. 2012) - - 26.25/0.737 24.65/0.653 25.03/0.629 22.17/0.615 23.45/0.741
Bilinear (Kirkland and Kirkland 2010) - - 27.55/0.788 25.42/0.679 25.54/0.646 22.69/0.635 24.21/0.767
Bicubic (Keys 1981) - - 28.42/0.810 26.00/0.702 25.96/0.667 23.14/0.657 24.91/0.787
ANR (Timofte, De Smet, and Van Gool 2013) 1.43MB 14.64 29.70/0.842 26.86/0.737 26.52/0.699 23.89/0.696 26.18/0.821
A+ (Timofte, De Smet, and Van Gool 2015) 15.17MB 155.34 30.27/0.860 27.30/0.750 26.73/0.709 24.33/0.719 26.91/0.848
CARN-M (Ahn, Kang, and Sohn 2018) 1.59MB 16.28 31.82/0.890 28.29/0.775 27.42/0.731 25.62/0.769 29.85/0.899
RRDB (Wang et al. 2018) 63.83MB 653.62 32.60/0.900 28.88/0.790 27.76/0.743 26.73/0.807 31.16/0.916
SRLUT (Jo and Kim 2021) 1.27MB 13.05 29.82/0.847 27.01/0.736 26.53/0.695 24.02/0.699 26.80/0.838
SPLUT-S (Ma et al. 2022) 5.5MB 56.32 30.01/0.852 27.20/0.743 26.68/0.702 24.13/0.706 27.00 0.843
SPLUT-M (Ma et al. 2022) 7MB 71.68 30.23/0.857 27.32/0.746 26.74/0.704 24.21/0.709 27.20 0.848
SPLUT-L (Ma et al. 2022) 18MB 184.32 30.52/0.863 27.54.0.752 26.87/0.709 24.46/0.719 27.70 0.858
MuLUT-SDY (Li et al. 2022) 3.82MB 39.15 30.40/0.860 27.48/0.751 26.79/0.709 24.31/0.714 27.52/0.855
MuLUT-SDY-X2 (Li et al. 2022) 4.06MB 41.60 30.60/0.865 27.60/0.754 26.86/0.711 24.46/0.719 27.90/0.863
@@
Experiments
Implementation Details
Datasets and Metrics. We train the proposed HKLUT on the DIV2K dataset (Agustsson and Timofte 2017), a popular dataset in the SR field. The DIV2K dataset provides 800 training images and 100 validation images with 2K resolution and the corresponding downsampled images. We use the widely adopted Peak Signal-to-Noise Ratio (PSNR) and structural similarity index (SSIM) (Wang et al. 2004) as evaluation metrics. Five well-known datasets: Set5 (Bevilacqua et al. 2012), Set14 (Zeyde, Elad, and Protter 2012), BSDS100 (Martin et al. 2001), Urban100 (Huang, Singh, and Ahuja 2015) and Manga109 (Matsui et al. 2017) are benchmarked. Following other LUT-based methods, we mainly focus on SR tasks with an upscaling factor of 4 (i.e., 4×4$\times$4 × SR). Besides, like MuLUT (Li et al. 2022), we measure and report the runtime, theoretical energy cost, and peak memory of generating a $1280\times 720$ HR image from a $640\times 360$ LR image.
Experimental setting. Before converting into LUTs, the network is trained with 800 training images in DIV2K for 200k iterations with a batch size of 16 on Nvidia RTX 3090 GPUs, each with a 24GB buffer. We utilize the Adam optimizer (Kingma and Ba 2014) ($\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=1\times 10^{-8}$) with the MSE loss to train the HKLUT. The initial learning rate is set to $5\times 10^{-4}$, which decays to one-tenth after 100k and 150k iterations, respectively. We randomly crop LR images into $48\times 48$ patches as input and enhance the dataset by random rotation and flipping. The runtime is measured on an Intel Core i5-10505 CPU with 16GB RAM, averaged over 10 runs.
Baselines. We evaluate HKLUT against several well-recognized SISR methods, such as interpolation-based methods (nearest neighbor (Bevilacqua et al. 2012), bilinear (Kirkland and Kirkland 2010) and bicubic (Keys 1981) interpolation), sparse-coding-based methods (ANR (Timofte, De Smet, and Van Gool 2013) and A+ (Timofte, De Smet, and Van Gool 2015)), DNN-based methods (CARN (Du et al. 2022) and RRDB (Wang et al. 2018)), and LUT-based methods (SRLUT (Jo and Kim 2021), SPLUT (Ma et al. 2022), and MuLUT (Li et al. 2022)).
Quantitative Results

### v2

@@ Abstract
Conventional super-resolution (SR) schemes make heavy use of convolutional neural networks (CNNs), which involve intensive multiply-accumulate (MAC) operations, and require specialized hardware such as graphics processing units. This contradicts the regime of edge AI that often runs on devices strained by power, computing, and storage resources. Such a challenge has motivated a series of lookup table (LUT)-based SR schemes that employ simple LUT readout and largely elude CNN computation. Nonetheless, the multi-megabyte LUTs in existing methods still prohibit on-chip storage and necessitate off-chip memory transport. This work tackles this storage hurdle and innovates hundred-kilobyte LUT (HKLUT) models amenable to on-chip cache. Utilizing an asymmetric two-branch multistage network coupled with a suite of specialized kernel patterns, HKLUT demonstrates an uncompromising performance and superior hardware efficiency over existing LUT schemes. Our implementation is publicly available at: https://github.com/jasonli0707/hklut.
1 Introduction
Single image super-resolution (SISR) is a long-standing problem in computer vision that involves generating a high-resolution (HR) image from a low-resolution (LR) image. The goal is to recover high-frequency details that are lost in a downsampled image. Classical non-deep-learning methods can be divided into two main categories: interpolation and sparse coding. Interpolation approaches Keys (1981); Kirkland and Kirkland (2010), such as bicubic interpolation, are simple and computationally efficient. They involve upsampling the LR image using a fixed kernel, e.g., a bicubic filter. However, they often suffer from limited generalizability and produce blurry images. Sparse-coding approaches Timofte et al. (2013, 2015), on the other hand, can achieve better results than the former but are computationally expensive and require hand-crafted features.
@@ 1 Introduction
Like many other problems in computer vision, deep learning has revolutionized SISR. The mapping between LR and HR images can be learned directly from data, often resulting in super-resolved images of much better quality than classical methods. However, deep neural networks (DNNs) require a large number of floating-point operations, resulting in a high power budget. Recently, researchers have explored new schemes that combine lookup tables (LUT) with deep-learning SISR Jo and Kim (2021); Li et al. (2022); Ma et al. (2022). Specifically, this approach pre-computes all possible input-output pairs of a deep SR network and stores them into an LUT for retrieval, thus skipping expensive computation during inference. As a result, LUT-based methods achieve inference speeds comparable to interpolation methods Jo and Kim (2021), but produce results with a learning-based quality. Moreover, LUT-based approaches are agnostic to software and hardware platforms, making them much more versatile than DNNs which require dedicated software and hardware support. Nonetheless, recent LUT-based SR schemes Li et al. (2022); Ma et al. (2022) mainly focus on improving performance by enlarging the receptive field (RF), while overlooking the memory constraints. This contradicts the trending edge AI that often runs on edge devices strained by limited resources, e.g., the popular Ultra96 FPGA board has $<1$MB on-chip memory. In fact, existing LUT-based schemes generally require multi-megabyte LUTs where off-chip storage is inevitable. In addition, existing LUT-based approaches may still require expensive interpolation and/or floating-point operations, resulting in higher peak memory and energy costs. This highlights the challenge as well as the need for significant LUT size reduction and better architecture design to address the storage and power issues, while upkeeping performance.
@@
To this end, this work systematically explores the relationship between the number of input pixels, RF, and SR performance. Our study leads to the design of a family of models called Hundred-Kilobyte Look-up Tables (HKLUTs) which offer a compelling balance between memory and performance. To eliminate the time-consuming and energy-intensive interpolation, a two-branch architecture as in SPLUT Ma et al. (2022) is employed. Yet instead of its symmetric branches Ma et al. (2022), we exploit the idea of effective receptive field (ERF) Luo et al. (2016) in different branches and reveal a larger RF is only required for the most-significant-bit (MSB) branch representing contextual semantic information, whereas a much smaller RF is sufficient for processing the least-significant-bit (LSB) branch.
@@
This key insight leads to an innovative asymmetric parallel structure, which largely reduces the storage space by nearly half without sacrificing performance. Next, we extend our model to multistage by cascading LUTs to enlarge the RF. Our proposed architecture, in contrast to prior works, allows information exchange between two branches during stage transition, resulting in an uncompromising performance. Moreover, we are the first to utilize progressive upsampling on LUT-based approaches, which further reduces the LUT size by half while improving the final performance. Compared to other state-of-the-art (SOTA) LUT-based schemes, HKLUTs deliver competitive performance while being >10×>10$\times$> 10 × smaller in size (cf. Fig. 1). The main contributions of this paper are threefold:
@@ 2 Related Work
The quest for high-quality SR on edge devices has fueled the interest and research of effective and efficient SISR schemes in recent years, broadly categorized into classical, deep learning and LUT-based Approaches.
@@
Classical Approaches. Basic linear interpolation methods Keys (1981); Kirkland and Kirkland (2010), though simple and efficient, are limited in their abilities to adapt to the image content and often result in blurry images. Early example-based work Timofte et al. (2013, 2015) learned sparse dictionaries over pixels or patches to build a compact representation between LR and HR pairs. During inference, the HR output is obtained using the pre-computed projection matrix with the input features. However, it remains time-consuming for sparse coding models to compute the sparse representations.
@@
Deep Learning Approaches. Deep learning methods have shown promising results in reconstructing image details for accurate super-resolution. Many deep learning SR works have been developed to reduce computational burdens by carefully adjusting the network structures, including, ESPCN Shi et al. (2016a), CARN Ahn et al. (2018), RRDB Wang et al. (2018) and IMDN Hui et al. (2019). LapSRN Lai et al. (2017) and its extension MS-LapSRN Lai et al. (2018) employ Laplacian Pyramid to generate multi-scale predictions and progressively reconstruct HR images in multiple steps. FMEN Du et al. (2022) shined in the NTIRE 2022 challenge with its low memory cost and short runtime through the enhanced residual block and high-frequency attention module. Although some models can be executed in real-time on GPUs, it is impractical to deploy them on edge devices with constrained hardware resources.
@@
LUT-based Approaches. While deep learning-based methods can produce impressive results, they often require heavy computations and large storage. To accelerate the inference of DNNs, dedicated software and hardware platforms such as CUDA and GPUs are required. This poses challenges for cost-sensitive and resource-restrictive edge devices such as smartphones and TVs. To address this, SRLUT Jo and Kim (2021) fuses DNN and LUT for SR applications by pre-computing and caching input-output pairs of a pre-trained SR network into a LUT for retrieval at test time. This approach eliminates expensive computation during DNN inference and achieves similar speed but better SR results than classical approaches. Two independent works around the same time, namely, MuLUT and SPLUT Li et al. (2022); Ma et al. (2022), propose to expand the RF covered by cascading LUTs. MuLUT proposes the use of multiple complementary LUTs together with the rotation ensemble to solve the intrinsic limitation of a small RF in SRLUT. However, MuLUT preserves the costly interpolation step of SRLUT. On the other hand, SPLUT Ma et al. (2022) exploits a novel parallel framework where the quantization loss is compensated by using two parallel branches that separately handle the most significant 4 bits and least significant 4 bits. This eliminates the time-consuming interpolation step for a faster inference speed. SPLUT also suggests enlarging the RF through feature aggregation along the horizontal and vertical directions. In addition, the Reconstructed Convolution (RC) module proposed in  Liu et al. (2023) can be transformed into 1D LUTs to efficiently expand the RF, thus improving performance. However, while these methods enhance performance compared to SRLUT, they require more storage space ($>1$MB), making them less suitable for edge devices.
3 Development of HKLUT
3.1 Reduced Number of Input Pixels
@@
The storage size of a LUT can be calculated using the formula $v^{n}\times r^{2}$, where $v$, $n$, and $r$ are the number of quantization levels, number of input pixels, and upscaling factor, respectively. The first term $v^{n}$ determines the number of entries, whereas the second term $r^{2}$ refers to the length of the output vector of each LUT. It is obvious that reducing the number of input pixels $n$ can result in an exponential reduction in storage space. Previous studies have primarily focused on expanding the RF by utilizing a fixed number of pixels (say, 4) to improve the final performance. SPLUT Ma et al. (2022) uses feature aggregation to expand the RF, namely, the output feature maps generated from the previous LUT are divided into groups along the channel dimension and then padded horizontally or vertically. The padded feature maps are then added to generate a single-channel feature map, which serves as the input for subsequent LUTs. However, in order to generate multi-channel feature maps, the size of intermediate LUTs grows linearly with the number of output channels. This goes against our primary goal, which is to reduce storage space. Subsequently, we deliberately omit feature aggregation. On the other hand, SRLUT Jo and Kim (2021) and MuLUT Li et al. (2022) employ rotation ensemble to increase the area being covered. The final prediction $\hat{y}_{i}$ is computed as follows:
@@
Figure 2: Three types of kernels covering a $3\times 3$ RF with rotation ensemble via different numbers of input pixels. (a) SRLUT (b) HDLUT (c) LLUT.
@@
3.2 Asymmetric Parallel Structure
Previous studies have attempted to reduce the size of LUTs by limiting the number of quantization levels to 17 and using interpolation to recover the quantization loss Jo and Kim (2021); Li et al. (2022). However, the interpolation stage is time-consuming and energy-intensive, hindering the practicality on edge devices. To address this, SPLUT Ma et al. (2022) proposes a parallel branch structure that avoids the need for interpolation. The structure divides the 8-bit input image into its most and least significant bits, viz. MSB and LSB branches each with 4 bits (viz. 16 levels) of information. The final SR image is obtained by fusing the results of the two branches at the last stage. While this approach gives promising results, it overlooks an important fact, namely, the effective receptive fields (ERFs) of MSB and LSB branches are completely different as they represent different kinds of information. Therefore, using the same sets of LUTs for both MSB and LSB branches might not be an optimal choice.
@@ 3.2 Asymmetric Parallel Structure
3.3 Multistage
LUT-based methods Ma et al. (2022); Li et al. (2022) have demonstrated the efficacy of cascaded stages to enlarge the RF for better performance. SPLUT Ma et al. (2022) splits the image into the MSB and LSB branches and deals with them separately with uniform quantization until the final addition, which prevents each branch from getting information from the other. MuLUT Li et al. (2022) utilizes quantization and interpolation to build the connection between stages, which requires later fine-tuning to compensate for the loss of information caused by quantization.
Here, we propose a new multistage framework as illustrated in Fig. 5. In each stage, we split the 8-bit input image into MSB and LSB branches for different LUT blocks to extract features and upscale the resolution. The features from both branches are then combined to generate the complete enlarged image. The input and output of each stage are 8-bit images with different resolutions, unless the upscaling factor is 1. This module can be stacked to create cascaded stages, where the latter stage can access information from the two branches of the former stages. Residual connections can be conveniently implemented. The entire flow does not require interpolation, which significantly speeds up inference.
@@ 3.3 Multistage
Nearest Bevilacqua et al. (2012) - - 26.25/0.737 24.65/0.653 25.03/0.629 22.17/0.615 23.45/0.741
Bilinear Kirkland and Kirkland (2010) - - 27.55/0.788 25.42/0.679 25.54/0.646 22.69/0.635 24.21/0.767
Bicubic Keys (1981) - - 28.42/0.810 26.00/0.702 25.96/0.667 23.14/0.657 24.91/0.787
ANR Timofte et al. (2013) 1.43MB 14.64 29.70/0.842 26.86/0.737 26.52/0.699 23.89/0.696 26.18/0.821
A+ Timofte et al. (2015) 15.17MB 155.34 30.27/0.860 27.30/0.750 26.73/0.709 24.33/0.719 26.91/0.848
CARN-M Ahn et al. (2018) 1.59MB 16.28 31.82/0.890 28.29/0.775 27.42/0.731 25.62/0.769 29.85/0.899
RRDB Wang et al. (2018) 63.83MB 653.62 32.60/0.900 28.88/0.790 27.76/0.743 26.73/0.807 31.16/0.916
SRLUT Jo and Kim (2021) 1.27MB 13.05 29.82/0.847 27.01/0.736 26.53/0.695 24.02/0.699 26.80/0.838
SPLUT-S Ma et al. (2022) 5.5MB 56.32 30.01/0.852 27.20/0.743 26.68/0.702 24.13/0.706 27.00 0.843
SPLUT-M Ma et al. (2022) 7MB 71.68 30.23/0.857 27.32/0.746 26.74/0.704 24.21/0.709 27.20 0.848
SPLUT-L Ma et al. (2022) 18MB 184.32 30.52/0.863 27.54.0.752 26.87/0.709 24.46/0.719 27.70 0.858
MuLUT-SDY Li et al. (2022) 3.82MB 39.15 30.40/0.860 27.48/0.751 26.79/0.709 24.31/0.714 27.52/0.855
MuLUT-SDY-X2 Li et al. (2022) 4.06MB 41.60 30.60/0.865 27.60/0.754 26.86/0.711 24.46/0.719 27.90/0.863
RCLUT Liu et al. (2023) 1.51MB 15.46 30.72/0.868 27.67/0.758 26.95/0.715 24.57/0.725 28.05/0.8655
@@
4 Experiments
4.1 Implementation Details
Datasets and Metrics. We train the proposed HKLUT on the DIV2K dataset Agustsson and Timofte (2017), a popular dataset in the SR field. The DIV2K dataset provides 800 training images and 100 validation images with 2K resolution and the corresponding downsampled images. We use the widely adopted Peak Signal-to-Noise Ratio (PSNR) and structural similarity index (SSIM) Wang et al. (2004) as evaluation metrics. Five well-known datasets: Set5 Bevilacqua et al. (2012), Set14 Zeyde et al. (2012), BSDS100 Martin et al. (2001), Urban100 Huang et al. (2015) and Manga109 Matsui et al. (2017) are benchmarked. Following other LUT-based methods, we mainly focus on SR tasks with an upscaling factor of 4. We also report the runtime, theoretical energy cost, and peak memory of generating a $1280\times 720$ HR image from a $640\times 360$ LR image.
Experimental setting. Before converting into LUTs, the network is trained with 800 training images in DIV2K for 200k iterations with a batch size of 16 on Nvidia RTX 3090 GPUs. We utilize the Adam optimizer Kingma and Ba (2014) ($\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=1e-8$) with the MSE loss to train the HKLUT. The initial learning rate is set to $5\times 10^{-4}$, which decays to one-tenth after 100k and 150k iterations, respectively. We randomly crop LR images into $48\times 48$ patches as input and enhance the dataset by random rotation and flipping. The runtime is measured on an Intel Core i5-10505 CPU with 16GB RAM, averaged over 10 runs.
Baselines. Aligning with prior work, we evaluate HKLUT against several well-recognized SISR methods, such as interpolation-based methods (nearest neighbor Bevilacqua et al. (2012), bilinear Kirkland and Kirkland (2010) and bicubic Keys (1981) interpolation), sparse-coding-based methods (ANR Timofte et al. (2013) and A+ Timofte et al. (2015)), DNN-based methods (CARN Du et al. (2022) and RRDB Wang et al. (2018)), and LUT-based methods (SRLUT Jo and Kim (2021), SPLUT Ma et al. (2022), MuLUT Li et al. (2022) and RCLUT Liu et al. (2023)).
4.2 Quantitative Results